The Elements of Thesis

A Thesis Presented to

The Academic Faculty

by

Perry H. Disdainful

In Partial Fulfillment of the Requirements for the Degree

Doctor of Philosophy

School of Electrical and Computer Engineering

Georgia Institute of Technology January 2001

The Elements of Thesis

Approved by:

Professor Ignatius Arrogant, Committee Chair

Professor Alpha Betic, Adviser

Professor Ivory Insular Professor General Reference (Another Department)

Professor Earl Grey

Professor Darjeeling Sweet

Date Approved

To myself, Perry H. Disdainful,

the only person worthy of my company.

PREFACE

Theses have elements. Isn’t that nice?

ACKNOWLEDGEMENTS

I want to “thank” my committee, without whose ridiculous demands, I would have graduated so, so, very much faster.

TABLE OF CONTENTS

DEDICATION iii PREFACE iv ACKNOWLEDGEMENTS v LIST OF TABLES vii LIST OF FIGURES viii SUMMARY ix I INTRODUCTION 1 II PREVIOUS WORK 2 III CONCLUSION 4 APPENDIX A — SOME ANCILLARY STUFF 5 REFERENCES 6 INDEX 7 VITA 8

LIST OF TABLES

1 A table, centered. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

LIST OF FIGURES

1 A first figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 A second figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

SUMMARY

Why should I provide a summary? Just read the thesis.

CHAPTER I INTRODUCTION

Every dissertation should have an introduction. You might not realize it, but the introduction should introduce the concepts, backgrouand, and goals of the dissertation.

Table 1: A table, centered.

Title  Author
War And Peace  Leo Tolstoy
The Great Gatsby F. Scott Fitzgerald

CHAPTER II PREVIOUS WORK

Some other research was once performed.

Figure 1: A first figure.

Figure 2: A second figure. 3

CHAPTER III CONCLUSION

APPENDIX A SOME ANCILLARY STUFF

Ancillary material should be put in appendices, which appear just before the bibliography.

REFERENCES

[1] Désarménien, J., “How to run TEX in french,” Tech. Rep. SATN-CS-1013, Computer Science Department, Stanford University, Stanford, California, Aug. 1984.

[2] Fuchs, D., “The format of TEX’s DVI files version 1,” TUGboat, vol. 2, pp. 12–16, July 1981.

[3] Fuchs, D., “Device independent file format,” TUGboat, vol. 3, pp. 14–19, Oct. 1982.

[4] Furuta, R. K. and MacKay, P. A., “Two TEX implementations for the IBM PC,” Dr. Dobb’s Journal, vol. 10, pp. 80–91, Sept. 1985.

[5] Knuth, D. E., “The WEB system for structured documentation, version 2.3,” Tech. Rep. STAN-CS-83-980, Computer Science Department, Stanford University, Stanford, California, Sept. 1983.

[6] Knuth, D. E., The TEX Book. Reading, Massachusetts: Addison-Wesley, 1984. Reprinted as Vol. A of Computers & Typesetting, 1986.

[7] Knuth, D. E., “Literate programming,” The Computer Journal, vol. 27, pp. 97–111, May 1984.

[8] Knuth, D. E., “A torture test for TEX, version 1.3,” Tech. Rep. STAN-CS-84-1027, Computer Science Department, Stanford University, Stanford, California, Nov. 1984.

[9] Knuth, D. E., TEX: The Program, vol. B of Computers & Typesetting. Reading, Massachusetts: Addison-Wesley, 1986.

[10] Lamport, L., L^{A}TEX: A Document Preparation System. User’s Guide and Reference Manual. Reading, Massachusetts: Addison-Wesley, 1986.

[11] Patashnik, O., BibTEXing. Computer Science Department, Stanford University, Stanford, California, Jan. 1988. Available in the BibTEX release.

[12] Patashnik, O., Designing BibTEX Styles. Computer Science Department, Stanford University, Jan. 1988.

[13] Samuel, A. L., “First grade TEX: A beginner’s TEX manual,” Tech. Rep. SATN-CS-83-985, Computer Science Department, Stanford University, Stanford, California, Nov. 1983.

[14] Spivak, M. D., The Joy of TEX. American Mathematical Society, 1985.

INDEX

7

VITA

Perry H. Disdainful was born in an insignificant town whose only claim to fame is that it produced such a fine specimen of a researcher.